OpenStudy (anonymous):

1/z^2-7z+6 - 1/z-1 = 1/2z-12

OpenStudy (anonymous):

It said the answer was 5 so I have no idea how they go to that conclusion

OpenStudy (anonymous):

\[ \frac{1}{z^2-7z+6}=\frac{1}{(z-1)(z-6)} \]

OpenStudy (anonymous):

Try multiplying both sides by \[ (z-1)(z-6) \]

OpenStudy (anonymous):

If I do that (z-1)(z-6) should cancel leaving only 1?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Just try it.

OpenStudy (anonymous):

With these equations, you want to get rid of the fractions.

OpenStudy (anonymous):

I keep on coming too 1-1z+6=1z-1

OpenStudy (anonymous):

\[ 1-(z-6)=\frac 12(z-1) \]

OpenStudy (anonymous):

Now solve for \(z\).